OpenStudy (anonymous):

How to factor 2x^2 - 1

pooja195 (pooja195):

its already factored .-.

OpenStudy (anonymous):

Ok so 2x^2 - 1 has solutions 2x^2 = 1 2x^2 = 1/2 sqr(x^2) = sqr(1/2) ok I am stuck here to finding the solutions

pooja195 (pooja195):

Thats not right .-. if you were to sqrt the 2 and then you'd be left with \[(\sqrt{2}+1)(\sqrt{2}-1)\] diffrence of squares .-.

OpenStudy (anonymous):

How do I get the solutions ?

pooja195 (pooja195):

Sqrt both numbers since its the difference of 2 squares ._.

pooja195 (pooja195):

a^2-b^2=(a+b)(a-b) a=sqrt(2)x b=1

OpenStudy (anonymous):

Ah ok I see now. Thank you.

pooja195 (pooja195):

\[(\sqrt{2}x+1)(\sqrt{2}x−1)\]

pooja195 (pooja195):

im sorry i meant that ^

OpenStudy (anonymous):

I got it. Thank you!

pooja195 (pooja195):

:) oh and this part too a^2-b^2 =\((\sqrt2 x)^2-1^2=2x^2-1\)

pooja195 (pooja195):

\(\huge\color{orange}{Welcome~~To~~OpenStudy~:)}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (anonymous):

TY